### v1

# Parity erasure: a foundational principle for indefinite causal order

Abstract
Processes with indefinite causal order can arise when quantum theory is locally valid. Here, we identify an information-theoretic principle, termed parity erasure, that completely characterizes such processes. Our characterization does not rely on the formalism of quantum theory itself, but instead is derived from a set of axioms for general operational probabilistic theories, and thus holds also for a large class of theories beyond quantum theory. This informational approach reveals a fundamental property of information exchange in scenarios with indefinite causal structure.
Introduction.— Over the past few decades, a fruitful line of research in the foundations of quantum mechanics has sought to reconstruct the quantum formalism from operational principles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. These reconstructions aim to interpret quantum theory not merely as a mathematical framework but as a set of fundamental properties of information processing. These operational principles have deepened our conceptual understanding of quantum mechanics and may serve as guiding constraints for potential extensions or modifications of the quantum formalism.
More recently, the scope of foundational investigations has expanded to scenarios in which the causal structure of quantum processes is dynamical, where it was shown that quantum theory accommodates more general processes with indefinite causal order (ICO) [13, 14, 15, 16, 17]. Such processes are described by the process matrix formalism, and can violate Bell-like inequalities for causal order [15, 18, 16, 19, 20], known as causal inequalities. They are expected to arise at the interface of quantum mechanics and general relativity [21, 22] and also shown to exist on time-delocalized quantum subsystems in standard quantum mechanics [23, 24].
The process matrix formalism is built on the concept of higher-order quantum operations, also known as quantum supermaps, which generalizes the quantum circuit model [14, 15]. Although supermaps can be defined within arbitrary general operational probabilistic theory (OPT) [1, 25, 26, 27, 6, 28, 29, 30, 31, 32] admitting the causality principle [27], processes with ICO remain largely abstract mathematical constructs lacking a clear operational interpretation. The possibility of ICO therefore opens a new frontier for the operational approach: if nature indeed allows such processes, it becomes crucial to understand whether the extension to ICO can be derived from operational principles without assumptions for higher-order operations.
In this paper, we identify a universal property of supermaps that holds in every causal OPT. Specifically, we show that supermaps satisfy a set of input–output constraints that we term parity erasure. Moreover, we prove that the parity erasure principle fully characterizes supermaps in every causal OPT which additionally satisfies the principle of local tomography and a property termed one-way-signaling decomposability. In particular, the process matrix formalism, which characterizes quantum supermaps, can be derived from the parity erasure principle. Our results thus provide an information-theoretic axiomatization of the process matrix formalism, revealing a fundamental constraint on information exchange within an indefinite causal structure.
Figure 1: Two-party supermaps. (a) Supermaps are transformations of channels; (b) A local input-output relation of a supermap is the statistics of an experiment in which the experimenter in each laboratory measures their incoming system and prepares an encoding of a random bit on the outgoing systems.
Supermaps in causal OPTs.— In a general OPT, the state space of a system $A$ is usually assumed to be a convex set ${\mathsf{St}}(A)$ which spans a finite-dimensional vector space ${\mathsf{St}}_{\mathbb{R}}(A)$. Every state in $\mathsf{St}(A)$ can be determined by the statistics of a choice of effects that span the dual space of ${\mathsf{St}}_{\mathbb{R}}(A)$. The composite system of $A$ and $B$ is denoted by $A\otimes B$. We assume the causality principle [27], which implies that for every system $A$ there is a unique deterministic effect that serves as the discarding operation $u^{A}$. The set of physical transformations from a system $A$ to a system $B$ is denoted by ${\mathsf{Transf}}(A\to B)$. The sets of deterministic states and deterministic transformations (also called channels) are denoted by ${\mathsf{St}}_{1}$ and ${\mathsf{Transf}}_{1}$, respectively. A test refers to a refinement of a channel, which corresponds to an instrument in quantum theory. States, effects, transformations and tests can be composed sequentially and in parallel to form a circuit. We refer to Refs. [27, 30, 32] for further details of the framework of OPTs.
Following the construction of the process matrix formalism [15], we consider a general OPT scenario [16] where there are $N$ separate experimenters, each of which can freely choose an operation in their laboratory. The operation of the $i$-th experimenter, denoted by $\left(\mathcal{M}_{a_{i}}^{(i)}\right)$, is a test consisting of probabilistic transformations from system $A_{i}$ to $X_{i}$ and producing an outcome $a_{i}$. The environment outside their laboratories is described by a supermap $\mathcal{S}$ transforming the local operations such that there is a valid joint probability distribution $p\left(\mathcal{M}_{a_{1}}^{(1)},\mathcal{M}_{a_{2}}^{(2)},\cdots\mid\mathcal{S}\right)$ for every choice of $\left(\mathcal{M}^{(i)}\right)_{i=1}^{N}$. Moreover, the map $\mathcal{S}$ is completely positive, that is, for every transformation $\widetilde{\mathcal{M}}^{(i)}$ from the composite system $A_{i}\otimes A^{\prime}_{i}$ to the composite system $X_{i}\otimes X^{\prime}_{i}$, the map $\mathcal{S}\otimes\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ produces a valid transformation from the composite system $A^{\prime}:=\bigotimes_{i=1}^{N}A^{\prime}_{i}$ to the composite system $X^{\prime}:=\bigotimes_{i=1}^{N}X^{\prime}_{i}$, where $\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ denotes the identity supermap on the set of transformations ${\mathsf{Transf}}(A^{\prime}\to X^{\prime})$. See Fig. 1(a) for an illustration in the case of two parties.
In addition, we assume the physicalization of readout [33], that is, every test can be realized as a channel followed by a measurement on an additional system, which indicates that supermaps can be completely characterized by their behaviors on channels.
The principle of parity erasure. — We now present a fundamental property of supermaps in their local input-output relations. A supermap is necessarily a channel sending the joint output of local parties to their joint input. This motivates us to study the statistics of experiments with the test in every local laboratory consisting of a measurement and a preparation, which in turn becomes a communication scenario through a supermap using local encoding and decoding strategies, as defined in the following and depicted in Fig. 1(b).
Definition 1
A local input-output relation of an $N$-partite channel $\mathcal{T}$ from $\bigotimes_{i=1}^{N}X_{i}$ to $\bigotimes_{i=1}^{N}A_{i}$ is a conditional joint probability distribution
$p(a_{1},\cdots,a_{N}\mid x_{1},\cdots,x_{N}):=\bigotimes_{i=1}^{N}e_{a_{i}}^{A_{i}}\left(\mathcal{T}\left(\bigotimes_{i=1}^{N}\rho^{X_{i}}_{x_{i}}\right)\right)$ (1)
where $x_{1},\cdots,x_{N}$ are random bits, $\rho^{X_{i}}_{x_{i}}$ refers to the state fed to the system $X_{i}$ according to the value of $x_{i}$, and $\left(e_{a_{i}}^{A_{i}}\right)_{a_{i}}$ refers to the effects of a measurement on the system $A_{i}$.
The causality principle imposes a basic constraint on the local input-output relations of supermaps, which forbids signaling from the future to the past in experiments in local laboratories. A similar constraint is studied in Ref. [34] in a framework beyond supermaps. Therefore, for every party $i$, the marginal probability distribution $p(a_{i}\mid x_{1},\cdots,x_{N})$ is independent of the value of $x_{i}$.
Another key observation is that a local input-output relation of a supermap can be regarded as a classical supermap [35]. In other words, the characterization of classical supermaps serves as a universal constraint on the local input-output relations of supermaps in any causal OPT.
In the two-party case, classical supermaps are known to be causally separable [15, 35], that is, every two-party classical supermap can be decomposed as a random mixture of ones compatible with fixed causal orders. When there is a fixed causal order between the two parties, one of the bits $x_{1}$ and $x_{2}$ cannot be decoded from the joint outcome $(a_{1},a_{2})$ in any local input-output relation, as causally-ordered classical supermaps can be realized by a circuit that discards the bit of the last party. Therefore, with a fixed causal order between the two parties or a random mixture of causal orders, the information about the parity $x_{1}\oplus x_{2}$ is not available from the joint measurement outcome $(a_{1},a_{2})$. It follows that the parity $x_{1}\oplus x_{2}$ cannot be predicted from the local input-output relations of any two-party supermap. This parity erasure property can be generalized to an arbitrary number of parties, which is formally stated in the following theorem.
Theorem 1 (Weak principle of parity erasure)
For any local input-output relation of a supermap in a causal OPT, and any nonempty subset $I$ of $\{1,\cdots,N\}$, the joint outcome $(a_{i})_{i\in I}$ is independent of the parity $\bigoplus_{i\in I}x_{i}$.
The proof of Theorem 1 uses the characterization of quantum supermaps in the process matrix formalism [36], which implies the constraint for arbitrary classical supermaps and thereby for all local input-output relations of supermaps in causal OPTs. The details of the proof are provided in Appendix A, which indicate a stronger version of parity erasure for quantum supermaps. In fact, for any causal OPT that satisfies local tomography (also known as local discriminability) [1], Theorem 1 is equivalent to a stronger version in which even joint measurements cannot gain information about the corresponding parity quantities. The stronger version of parity erasure is formally stated with the following definition and corollary.
Definition 2
An $N$-partite channel $\mathcal{T}$ from $\bigotimes_{i=1}^{N}X_{i}$ to $\bigotimes_{i=1}^{N}A_{i}$ is parity-erasure if for any choice of local encoding strategies $\left(\rho_{x_{i}}^{X_{i}}\right)$ and any nonempty subset $I$ of $\{1,\cdots,N\}$, the marginal state $\bigotimes_{i\notin I}u^{A_{i}}\left(\mathcal{T}\left(\bigotimes_{i=1}^{N}\rho_{x_{i}}^{X_{i}}\right)\right)$, obtained by discarding the systems $(A_{i})_{i\notin I}$, is independent of the parity $\bigoplus_{i\in I}x_{i}$.
Corollary 1 (Strong principle of parity erasure)
In a causal OPT that satisfies the principle of local tomography, supermaps are parity-erasure channels.
Characterizing supermaps from principles.— We have seen that every supermap can be described by a channel whose local input-output relations satisfy the parity erasure principle. We now show that the converse is also true in a class of theories including quantum theory, that is, in these theories, every parity-erasure channel corresponds to a supermap.
We assume that local tomography and the following property of one-way-signaling decomposability are satisfied.
Figure 2: One-way-signaling decomposability. Under this assumption, every transformation satisfying the no-influence condition $X_{1}\not\to A_{1}$ can be implemented by a circuit with a slot from $A_{1}$ to $X_{1}$.
Axiom 1 (One-way-signaling decomposability)
Every bipartite channel $\mathcal{U}\in{\mathsf{Transf}}_{1}(X_{1}\otimes X_{2}\to A_{1}\otimes A_{2})$ that satisfies the no-influence condition $X_{1}\not\to A_{1}$ (that is, the marginal output on the subsystem $A_{1}$ does not depend on the marginal input on the subsystem $X_{1}$) can be implemented by concatenating a transformation $\mathcal{E}\in{\mathsf{Transf}}_{1}(X_{2}\to A_{1}\otimes M)$ and another transformation $\mathcal{D}\in{\mathsf{Transf}}_{1}(X_{1}\otimes M\to A_{2})$ through a system $M$, as visualized in Fig. 2.
For classical channels, one-way-signaling decomposability follows directly from the chain rule of probability theory, and this property has also been established for quantum channels [37]. This property guarantees that any channel for which a given input subsystem cannot influence a given output subsystem admits an explicit circuit realization in which an arbitrary transformation between the two subsystems can be probed.
With local tomography and one-way-signaling decomposability, it follows that any $N$-partite parity-erasure channel can act as a supermap through a procedure in which the local transformations of the $N$ parties are inserted one-by-one into a sequence of decompositions of the parity-erasure channel. The proof is established by a recursive property of extended parity-erasure channels. Extended $k$-partite parity-erasure channels from $X_{1}\otimes\cdots\otimes X_{k}\otimes P$ to $A_{1}\otimes\cdots\otimes A_{k}\otimes F$ refer to those that are reduced to parity-erasure channels after preparing an arbitrary state on the system $P$ and discarding the system $F$. Such channels admit a one-way-signaling decomposition with a slot from $A_{i}$ to $X_{i}$ ($i\in\{1,\cdots,k\}$) because the parity-erasure property for a single party reduces to the no-influence condition $X_{i}\not\to A_{i}$. The recursive property is stated in the following lemma (see the proof in Appendix B).
Lemma 1
Let $\mathcal{T}$ be an extended $k$-partite parity-erasure channel from $X_{1}\otimes\cdots\otimes X_{k}\otimes P$ to $A_{1}\otimes\cdots\otimes A_{k}\otimes F$, and let $\mathcal{C}\in{\mathsf{Transf}}_{1}(A_{k}\otimes P^{\prime}\to X_{k}\otimes F^{\prime})$. The transformation obtained by inserting $\mathcal{C}$ into the one-way-signaling decomposition of $\mathcal{T}$ is an extended (k1)(k-1)-partite parity-erasure channel from $X_{1}\otimes\cdots\otimes X_{k-1}\otimes(P\otimes P^{\prime})$ to $A_{1}\otimes\cdots\otimes A_{k-1}\otimes(F\otimes F^{\prime})$.
The characterization of supermaps from principles is summarized in the following theorem.
Theorem 2
In a causal OPT that satisfies local tomography and one-way-signaling decomposability, an $N$-partite channel corresponds to an $N$-party supermap if and only if it is parity-erasure. In other words, supermaps in such theories are completely characterized by the parity erasure principle.
Discussion and outlook.— We have provided an information-theoretic characterization of processes with ICO by identifying the parity erasure principle that is universally satisfied by supermaps in all causal OPTs. Under the minimal assumptions of local tomography and one-way-signaling decomposability, this principle determines the admissible set of supermaps and, when applied to quantum theory, yields precisely the process matrix formalism. Our results place ICO within the operational reconstruction program and show that it can be seen as arising from a fundamental property of information exchange rather than from purely formal extensions of the circuit model.
An important direction for future work is to explore the implications of the parity erasure principle for concrete applications of ICO, such as the characterization of correlations with ICO [38, 39, 40] and cyclic quantum causal models [41, 42].
From a broader perspective, our results suggest that information-theoretic principles may play a central role in any theory with dynamical causal structure, including prospective theories of quantum gravity. They also open the door to systematic investigations of ICO in post-quantum theories and to a clearer identification of what advantages may be expected in such processes.
Acknowledgment.— This work was supported by the F.R.S.-FNRS under project CHEQS within the Excellence of Science (EOS) program, by the ID# 62312 grant from the John Templeton Foundation, as part of the “The Quantum Information Structure of Spacetime” Project (QISS), and by the ID# 63683 grant from the John Templeton Foundation, as part of the “WithOut SpaceTime” Project (WOST). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. O. O. is a Senior Research Associate of the Fonds de la Recherche Scientifique (F.R.S.–FNRS).
Appendix A Proof of Theorem 1
An $N$-party quantum supermap $\mathcal{S}$ can be characterized by the following conditions on its Choi representation $S$ [36]:
$S\geq 0\,$, (2)
$\operatorname{Tr}(S)=\prod_{i=1}^{N}d_{X_{i}}\,$, (3)
$\forall~\text{nonempty}~I\subseteq\{1,\cdots,N\}$
${}_{\prod_{i\in I}(1-X_{i})\prod_{i\notin I}A_{i}X_{i}}S=0\,$, (4)
where we have used the notation $_{[X]}S:=\operatorname{Tr}_{X}[S]\otimes\frac{I_{X}}{d_{X}}$ for a quantum system $X$ of dimension $d_{X}$, and the shorthand $_{[\sum\alpha_{X}X]}S:=\sum_{X}\alpha_{X}\cdot{}_{[X]}S$ for a linear combination of the terms $\{{}_{[X]}S\}_{X}$ associated with a collection of systems indexed by $X$ (for example, $_{[(1-X)Y]}S$ serves as a shorthand for $_{[Y]}S-{}_{[XY]}S$).
To analyze local input-output relations, let the systems $X_{1},X_{2},\cdots,X_{N}$ be classical bits. $S$ can be viewed as a classical-to-quantum channel which outputs a quantum state $\Psi_{\vec{x}}$ when the input are classical bits $\vec{x}=(x_{1},\cdots,x_{N})$, i.e.,
$S=\sum_{\vec{x}\in\{0,1\}^{N}}|\vec{x}\rangle\langle\vec{x}|\otimes\Psi_{\vec{x}}\,.$ (5)
Conditions (2) and (3) are necessary for $S$ to represent a quantum channel. Conditions (4) are equivalent to the following equations:
$\forall~\text{nonempty}~I\subseteq\{1,\cdots,N\}$
$\bigotimes_{i\notin I}\operatorname{Tr}_{A_{i}}\left(\sum_{\vec{x}}\prod_{i\in I}(-1)^{x_{i}}\Psi_{\vec{x}}\right)=0\,.$ (6)
This proves that any quantum supermap is parity-erasure (see Definition 2). Since classical supermaps can be viewed as a special case of quantum supermaps [35], this implies that classical supermaps are also parity-erasure. Since the local input-output relations for a supermap in any causal OPT correspond to classical supermaps, this proves Theorem 1.
Appendix B Proof of Lemma 1
For brevity, we introduce the following shorthand notations
$u^{\overline{I}}$ $:=\bigotimes_{i\notin I}u^{A_{i}}\,$, (7)
$\rho^{\overline{I}}$ $:=\bigotimes_{i\notin I}\rho^{X_{i}}_{i}\,.$ (8)
Let $\mathcal{T}_{\mathcal{C}}$ denote the channel obtained by inserting $\mathcal{C}$ into the one-way-signaling decomposition of $\mathcal{T}$. It suffices to show that for any nonempty subset $I\subseteq\{1,\cdots,k-1\}$, and for arbitrary states on the corresponding systems,
$u^{\overline{I}}u^{F}u^{F^{\prime}}\mathcal{T}_{\mathcal{C}}\left(\bigotimes_{i\in I}\left(\rho^{X_{i}}_{i,0}-\rho^{X_{i}}_{i,1}\right)\otimes\rho^{\overline{I}}\otimes\sigma^{P}\otimes\xi^{P^{\prime}}\right)=0\,.$ (9)
Let $\mathcal{T}_{\sigma}$ denote the channel obtained by preparing $\sigma$ on the system $P$ and discarding the system $F$. And let $\mathcal{C}_{\xi}\in{\mathsf{Transf}}_{1}(A_{k}\to X_{k})$ denote the channel obtained by preparing $\xi$ on the system $P^{\prime}$ and discarding the system $F^{\prime}$. Since local tomography allows for a local characterization of transformations [27], $\mathcal{C}_{\xi}$ can be written as a linear combination of the parallel composition of an effect on $A_{k}$ and a state on $X_{k}$, that is,
$\mathcal{C}_{\xi}(\cdot)=\sum_{j}c_{j}e_{j}^{A_{k}}(\cdot)\rho_{k,j}^{X_{k}}\,$, (10)
with real coefficients $(c_{j})$ and states $\rho_{k,j}^{X_{k}}\in{\mathsf{St}}_{1}(X_{k})$. It follows that $\mathcal{C}_{\xi}$ can be written as
$\mathcal{C}_{\xi}(\cdot)=u^{A_{k}}(\cdot)\pi^{X_{k}}+\sum_{j}c_{j}e_{j}^{A_{k}}(\cdot)\left(\rho_{k,j}^{X_{k}}-\pi^{X_{k}}\right)\,$, (11)
where $\pi^{X_{k}}\in{\mathsf{St}}_{1}(X_{k})$ is an arbitrary state. To prove Eq. (9), it suffices to show that
$u^{\overline{I}}u^{A_{k}}\mathcal{T}_{\sigma}\left(\bigotimes_{i\in I}\left(\rho^{X_{i}}_{i,0}-\rho^{X_{i}}_{i,1}\right)\otimes\rho^{\overline{I}}\otimes\pi^{X_{k}}\right)=0$ (12)
and
$u^{\overline{I}}\mathcal{T}_{\sigma}\left(\bigotimes_{i\in I}\left(\rho^{X_{i}}_{i,0}-\rho^{X_{i}}_{i,1}\right)\otimes\rho^{\overline{I}}\otimes\left(\rho^{X_{k}}_{k,j}-\pi^{X_{k}}\right)\right)=0\,.$ (13)
Both Eq. (12) and Eq. (13) follow from the fact that $\mathcal{T}_{\sigma}$ is parity-erasure. This completes the proof.